4) You do 200 J of work in 61 seconds, and your friend does the same work in 1 minute. It follows that your friend is more powerful than you are(T,F)
Answer #1

W = 200 Joule

t1 = 61s and t2 = 1mint = 60s

P1 = 200/61 = 3.278

P2 = 200/60 = 3.334

=) P2 > P1

That means my friend is more powerful than me. Hence statement is true .